If y varies inversely as X^2, and y=-8, when x=2, find y if x=3.
Any help would be great! Thank you

Answer by stanbon(57282)

If y varies inversely as X^2, and y=-8, when x=2, find y if x=3.
---------------
y=k/x^2 since y varies inversely as x^2.
Need to find "k"
8=k/2^2
k=32
EQUATION:
y=32/x^2
If x=3 then:
y=32/3^2=32/9
y=3 5/9
Cheers,
Stan H.